OpenStudy (anonymous):

Solve each problem by using the quadratic formula. Write solutions in simplest radical form. 2x^2 + 2x- 1 = 0

OpenStudy (anonymous):

so ax^2 + bx + c = 0

OpenStudy (anonymous):

\[(-b \pm \sqrt{b^2 - 4 a c}) /2a\]

OpenStudy (anonymous):

just plug in the values your equation is 2x^2 + 2x - 1 = 0 ax^2 + bx + c = 0

OpenStudy (anonymous):

tell me what a = , b = , and c =

OpenStudy (anonymous):

are they a=2 b=2 c= -1 ?

OpenStudy (anonymous):

perfect now plug them into the quadratic equaiton

OpenStudy (anonymous):

and you must rememeber the pm

OpenStudy (anonymous):

pm -> plus and minus

OpenStudy (anonymous):

because you will have two solutions one for a positive square root, and one for a negative square root

OpenStudy (anonymous):

\[-2\pm \sqrt{12} /4\] this is what im getting so far am i on the right track

OpenStudy (anonymous):

yeah. see that plus and minus you have to look at it like this -2 + sqrt(12)/4 and -2 - sqrt(12)/4

OpenStudy (anonymous):

so you will get two answers

OpenStudy (anonymous):

would the first answer be 1 and the second answer be -2

OpenStudy (anonymous):

you will get -1 +- sqrt(3)/2

OpenStudy (anonymous):

i have to go

OpenStudy (anonymous):

you will get -1 +- sqrt(3)/2